OpenStudy (anonymous):

Lim x goes to 1 x^2-1/|x-1| anyone can find this limit?

OpenStudy (anonymous):

\[\lim_{x\to 1}\frac{x^2-1}{|x-1|}\]

OpenStudy (anonymous):

seems unlikely since this is is a piecewise function it is one thing if \(x>1\) and quite another if \(x<1\)

OpenStudy (anonymous):

if \(x>1\) then \(|x-1|=x-1\) when you factor and cancel you get \[x+1\]

OpenStudy (anonymous):

you get something else if \(x<1\) try it and see

OpenStudy (anonymous):

I wonder if I should use left and right limit to do this

OpenStudy (anonymous):

you have no choice but to do that, since \(|x-1|\) is a piecewise function which changes definition at \(x=1\)

OpenStudy (anonymous):

lets take it step by step

OpenStudy (anonymous):

if \(x>1\) then \(|x-1|=x-1\) right?

OpenStudy (anonymous):

did i lose you there?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

if \(x>1\) then \(|x-1|=x-1\)so your function is \[\frac{x^2-1}{x-1}=\frac{(x+1)(x-1)}{x-1}=x+1\] making the limit as \(x\to 1^+=1+1=2\)

OpenStudy (anonymous):

I see, now the left limit I guess

OpenStudy (anonymous):

left limit is different because if \(x<1\) then \(|x-1|=1-x\)

OpenStudy (anonymous):

so the limit doesnt exist right

OpenStudy (anonymous):

no the two sided limit does not exist

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

Thank you